The set X = {12, 15, a, b} has a mean 15.5 and median 16. It is given that set X is arranged in ascending order.

Determine the values of a and b. [2 marks]

Find the standard deviation of set X. [1 mark]

When each number in set X is decreased by 10, find the mean and standard deviation. [2 marks]

If each number of X is increased by 100%, find the mean and standard deviation.


1
Expert's answer
2021-05-11T15:14:18-0400

The given set is-


X={12,15,a,b}


"\text{Mean} =\dfrac{12+15+a+b}{4}

\\[9pt]

15.5=\dfrac{27+a+b}{4}

\\[9pt]

\Rightarrow a+b=15.5\times 4-27

\\[9pt]

\Rightarrow a+b=35~~~~~-(1)"


"Median =(\dfrac{N}{2}+1)^{th} term


\\[9pt]
16=(\dfrac{4}{2}+1)^{th} term
\\[9pt]


16=3^{th} term"


hence a"=16" and "b=35-16=19"


Now The set X is {12,15,16,19}


Variance "=\dfrac{\sum(x_i-\bar{x})^2}{4}=\dfrac{(-3.5)^2+(-0.5)^2+(0.5)^2+(3.5)^2}{4}=\dfrac{2405.5}{4}=612.625"


Standard deviation "=\sqrt{\text{variance}}=\sqrt{612.625}=24.75"



When each number is decreased by 10 , The new set became {2,5,6,9}


New Mean "=\dfrac{2+5+6+9}{4}=\dfrac{22}{4}=5.5"


Standard deviation remains same i.e. 24.75


When each number is increased by 100%, The new set is {24,30,32,38}


New mean "=\dfrac{24+30+32+38}{4}=\dfrac{124}{4}=31"


Standard deviation remains same as 24.75